Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Outdoor National Championships

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Fassaroe Shooting Club
Date July 7th, 8th, 2001
Discipline 50m Three-Position and 50m Prone Rifle

DURC Attendees

Mark Dennehy  John Keeney
Alan Smyth  Adrian Costigan
Peter Byrne  Conor McDermottroe

Joe Lalor

Scores Three Position Rifle

Mark Dennehy 937(ex1200)

Prone Rifle

John Keeney 577
Alan Smyth 569
Ray Reilly 566
Conor McDermottroe 562
Mark Dennehy 561
Peter Byrne 557
Adrian Costigan 554
Joe Lalor 522

Prizes Won

John Keeney First in Class B, Silver Medal overall
Alan Smyth First in Class D
Adrian Costigan Second in Class C
Peter Byrne Third in Class D
Joe Lalor Third in Class C

Report

DURC entered a reasonable contingent in this year’s 50m Prone Nationals this weekend and all did extremely well. John Keeney managed to score 577 and take the overall second place, making up for Ray Reilly losing the title of National Champion somewhat. Alan Smyth also shot well, making it to the 10-shot final and placing 7th to do so. Newcomers to the 50m sport did well also, with Adrian Costigan, Joe Lalor, and Peter Byrne walking away with prizes, though Conor McDermottroe was only denied a prize by a contraversial classification decision.

Only one competitor entered the three-position match however, and I did not shoot as well as I had hoped, coming in fourth but increasing my score by about a hundred points from my last national 3P match. DURC will begin training for 3P in earnest this coming academic year using Air Rifles due to range restictions, so hopefully we will have a better team fielded next year for the 3P match.

In summary then, DURC has lost the Gold medal in 50m Prone, but has taken the silver by way of compensation; Five of the eight shooters who competed took prizes, and two made it to the final; and we shared joint first for the highest-scoring team at the competition with Fassaroe Sporting Club (who can count two National Champions and an Olympian amongst their team against our beginners and two or three “old hands” with a few years behind them). Proving once again that DURC has fully recovered from the 1996-98 slump to take it’s place as one of the best clubs in Ireland.

Prone Rifle Competition Winners : John Keeney (Silver), Alan Lewis (Gold) and Rhona Barry (Bronze)

Mark Dennehy shooting Prone and Standing at the 3P event John Keeney shooting at the Prone Rifle Championships and Alan Lewis shooting kneeling at the 3P competition Rhona Barry (Atlanta Olympics) and Alan Lewis (Sydney Olympics) shooting Prone during the 3P championships

Misé le meas,

Mark Dennehy, DURC Secretary